OpenStudy (fran345):

In physics, if a moving object has a starting position at s 0, an initial velocity of v 0, and a constant acceleration a, then the position S at any time t > 0 is given by: S = at 2 + v 0 t + s 0. Solve for the acceleration, a, in terms of the other variables. For this assessment item, you can use ^ to show exponents and type your answer in the answer box, or you may choose to write your answer on paper and upload it.

OpenStudy (sooobored):

shouldnt it be \[\frac{1}{2} a t^2\]

OpenStudy (fran345):

yeah, I forgot to add that

OpenStudy (sooobored):

well, the equation is already given, s0 is given v0 is given a great start would be substituting those value into the equation

OpenStudy (sooobored):

so the resulting equation would be?

OpenStudy (fran345):

So it's S=1/2at^2+v0+s0?

OpenStudy (sooobored):

no, s0= 0 and v0 = 0, as stated in your problem

OpenStudy (fran345):

so I drop the v and s? I'm confused?

OpenStudy (sooobored):

so you substitute the value zero, into the equation for v0 and s0 \[S = \frac{1}{2}at ^2 + 0 t +0.\]

OpenStudy (sooobored):

and since zero is zero, you can just ignore them and you get the equation \[S=\frac{1}{2}at^2\]

OpenStudy (sooobored):

now since the problem just wants you to solve for "a" in terms of the other variables

OpenStudy (fran345):

would it be a=2s/t^2